OpenStudy (anonymous):

prove that if x1,x2,...xn are n real numbers in the closed interval [a,b], then (x1+x2+...+xn)/n is in the closed interval [a,b]

OpenStudy (anonymous):

this is real analysis problem. any ideas?

OpenStudy (anonymous):

i know need to use math induction... getting stuck at the end.... missing something

myininaya (myininaya):

go joe!

OpenStudy (anonymous):

I would start by noting that if all n numbers were "a", then we would have the sum: \[a+a+a+\cdots+a = na\]This is also the smallest sum we could have. Likewise, if all the numbers were b, we would have: \[b+b+b+\cdots +b = nb\]which is the largest sum we could possible have. So: \[a+a+a+\cdots +a\leq x_1+x_2+\cdots +x_n\leq b+b+b+\cdots +b\] \[\iff na\leq x_1+x_2+\cdots +x_n \leq nb\] \[\iff a\leq \frac{x_1+x_2+\cdots+x_n}{n}\leq b\]

OpenStudy (zarkon):

\[\min(\{x_1,x_2,\ldots,x_n\})\leq\bar{x}\leq\max(\{x_1,x_2,\ldots,x_n\})\]

myininaya (myininaya):

joe thats cute i like it

OpenStudy (anonymous):

lolol :) zarkons is so much better looking though (as always :P)

myininaya (myininaya):

i like yours better no offense zarkon

myininaya (myininaya):

its remedial is why i like it

OpenStudy (zarkon):

nothing wrong with what you have...I just have a tighter bound

OpenStudy (zarkon):

really the same Idea

myininaya (myininaya):

like zarkon uses doctor notation

myininaya (myininaya):

he thinks he better than all of us non-doctors

myininaya (myininaya):

jk

OpenStudy (anonymous):

yeah. @myin for some reason that sounds like an insult :P lolol i know its not, just saying

OpenStudy (zarkon):

lol

myininaya (myininaya):

no its not an insult

OpenStudy (anonymous):

"down to earth" language? lolol

myininaya (myininaya):

i was just trying to say you broke it all the way down

myininaya (myininaya):

totally easy to understand

myininaya (myininaya):

i mean zarkon's was easy to understand too

myininaya (myininaya):

just ignore me

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

thank you, guys! awesome.... i suppose to use induction method, though... I got p(1)- true; p(k) - assumed to be true. then need to proof: a<(x1+x2+...xk+x(k+1))/(k+1) <b

OpenStudy (anonymous):

oh its required to prove by induction? eek

OpenStudy (zarkon):

it really is the same idea...replace a in joe's proof with \[x_{(1)}\] and b with \[x_{(n)}\] then proceed with the rest of his proof.

OpenStudy (anonymous):

close, but not the same

OpenStudy (anonymous):

i got: a<(x1+x2+...x(k+1))/(k+1)<b a+a/k<p(k) + x(k+1)/(k+1) <b+b/k x(k+1)= means index not multipl

OpenStudy (anonymous):

not to worry... I got it! THANK YOU EVERYBODY!

OpenStudy (anonymous):

im posting my inductive proof anywhos <.< i spent some time writing it out lolol

OpenStudy (anonymous):

Skipped the base step, just showed the inductive step

OpenStudy (zarkon):

you posted the same thing twice...missing last page?

OpenStudy (anonymous):

whoops! where'd it go....

OpenStudy (anonymous):

my bad >.<

OpenStudy (zarkon):

if you add a Basis Step then it will be perfect ;)

myininaya (myininaya):

a doctor wanting to add a basic step i don't know sounds fishy

OpenStudy (anonymous):

yeah, if this was my homework it would totally be there. Most times when people need help with induction, its the inductive step thats killing them, so ive gotten used to skipping to it >.< lolol

myininaya (myininaya):

that means its not really basic

OpenStudy (zarkon):

nice work Joe!

myininaya (myininaya):

yes joe is awesome

myininaya (myininaya):

and so is zarkon

OpenStudy (anonymous):

thank you :) you guys are the awesome ones <.<

myininaya (myininaya):

i'm going to sleep

OpenStudy (zarkon):

sounds like a good idea...I started back to work this week

myininaya (myininaya):

i missed a cool trig problem below this thread :(

OpenStudy (anonymous):

i have class tomorrow morning >.< and some people want to meet early and get Number Theory/Problem SOlving work done >.<

OpenStudy (zarkon):

I don't teach on Thursdays :)

myininaya (myininaya):

i was too busy picking on joe and i didn't realize

OpenStudy (anonymous):

lucky o.O

OpenStudy (anonymous):

lol <.<

myininaya (myininaya):

ok later

OpenStudy (zarkon):

later

OpenStudy (anonymous):

lates, im out as well.

OpenStudy (zarkon):

good night Joe/myininaya